# Thermal one-loop self-energy correction for hydrogen-like systems: relativistic approach

### v1

@@ IV Conclusions
Previously, in the work [Lopez_2025], it was shown that this QED correction, generalized to the case of finite temperature, includes thermal corrections known from quantum mechanical theory (Stark and Zeeman effects, higher multi-pole polarizabilities). Moreover, a careful transition to the non-relativistic limit, which restores the specified effects, made it possible to point out that in the thermal one-loop correction there are additional contributions related to relativistic effects that were not previously taken into account. In the same work [Lopez_2025], it was stated that the thermal self-energy correction of a bound electron can be accurately calculated within a fully relativistic approach; see also [Porsev]. Here is the corresponding analysis using partial decomposition by spherical functions for the thermal self-energy operator.
@@
As the first result of a fully relativistic approach, we’ve calculated the thermal shifts of the ground state for hydrogen and hydrogen-like ions with an arbitrary nuclear charge $Z$. Specifically, from Table 1, it follows that the calculation method is most accurate for light ions (with a small $Z$). Another consequence is the behavior of the thermal shift with respect to $Z$. The decrease as $1/Z^{4}$ has been verified with high accuracy, see Fig. 2. The dependence $1/Z^{4}$ was obtained for the non-relativistic limit and dipole approximation [Lopez_2025]. Therefore, a slight deviation for large $Z$ on the graph in Fig. 2 clearly demonstrates the importance of the relativistic approach in highly charged (H-like in particular) ions.
@@
Based on the analysis above, we have revised the quantum mechanical values of thermal level shifts in a hydrogen atom [farley], see Table 2. Deviation from generally accepted values (see corresponding sub-strings in Table 2) is explained by several reasons. First and foremost, the need to account for the Lamb shift [PRA_2015, jetp2022] and the fine structure of atomic levels [reex]. The next reason is the approximation for frequency integration, presented in [farley] and still frequently used today for complex atomic systems [Glukhov_2016, PhysRevA.110.043108]. We emphasize that this work presents fully relativistic calculations of the expressions representing the one-loop thermal shift, which is unequivocally consolidated with quantum mechanical theory [Lopez_2025].
@@
The achieved precision in calculating thermal shifts of atomic energy levels has made it possible to accurately determine the frequency shift for transitions between fine and hyperfine sub-levels at various temperatures and for various hydrogen-like systems, see Tables 3, 5 and 6. The obtained values were used to study the behavior of shifts in hyperfine intervals as a function of both temperature and $Z$, see Figs. 5, 6.
@@
The most important result of the research presented in this work is the discovery of the magic temperature at which the thermal frequency shift is zero [PhysRevA.100.023417]. For the most accurately measured frequencies in the hydrogen atom, as well as for some hydrogen-like ions, the results are presented in Table 4, see also Figs. 3, 4. It should be noted that there is a significant difference in the magic temperature for various hydrogen-like systems. For example, for doubly ionized lithium, the thermal shift of the two-photon 1s3s1s-3s transition frequency vanishes at a temperature significantly exceeding room temperature. At the same time, for lighter systems, magic wavelengths appear achievable in experiments involving two-photon transitions 1s2s1s-2s, 1s3s1s-3s (except for the helium ion in the latter case). It should also be noted that the analysis, taking into account the hyperfine structure of atomic levels, revealed that the latter is insignificant for determining $T_{m}$. This can be seen directly from Table 6: the thermal shifts to the hyperfine intervals of the ground and $2s$ states were always positive, differing by approximately an order of magnitude and at the same time negligibly small. For a hydrogen atom and the most accurately measured transition frequencies of 1s2s1s-2s and 1s3s1s-3s [CODATA-2021], the values obtained are $T_{m}=65.9$ K and $T_{m}=29.7$ K, respectively.

### v2

@@ IV Conclusions
Previously, in the work [26], it was shown that this QED correction, generalized to the case of finite temperature, includes thermal corrections known from quantum mechanical theory (Stark and Zeeman effects, higher multi-pole polarizabilities). Moreover, a careful transition to the non-relativistic limit, which restores the specified effects, made it possible to point out that in the thermal one-loop correction there are additional contributions related to relativistic effects that were not previously taken into account. In the same work [26], it was stated that the thermal self-energy correction of a bound electron can be accurately calculated within a fully relativistic approach; see also [31]. Here is the corresponding analysis using partial decomposition by spherical functions for the thermal self-energy operator.
@@
As the first result of a fully relativistic approach, we’ve calculated the thermal shifts of the ground state for hydrogen and hydrogen-like ions with an arbitrary nuclear charge $Z$. Specifically, from Table 1, it follows that the calculation method is most accurate for light ions (with a small $Z$). Another consequence is the behavior of the thermal shift with respect to $Z$. The decrease as $1/Z^{4}$ has been verified with high accuracy, see Fig. 2. The dependence $1/Z^{4}$ was obtained for the non-relativistic limit and dipole approximation [26]. Therefore, a slight deviation for large $Z$ on the graph in Fig. 2 clearly demonstrates the importance of the relativistic approach in highly charged (H-like in particular) ions.
@@
Based on the analysis above, we have revised the quantum mechanical values of thermal level shifts in a hydrogen atom [14], see Table 2. Deviation from generally accepted values (see corresponding sub-strings in Table 2) is explained by several reasons. First and foremost, the need to account for the Lamb shift [42, 50] and the fine structure of atomic levels [21]. The next reason is the approximation for frequency integration, presented in [14] and still frequently used today for complex atomic systems [16, 45]. We emphasize that this work presents fully relativistic calculations of the expressions representing the one-loop thermal shift, which is unequivocally consolidated with quantum mechanical theory [26].
@@
The achieved precision in calculating thermal shifts of atomic energy levels has made it possible to accurately determine the frequency shift for transitions between fine and hyperfine sub-levels at various temperatures and for various hydrogen-like systems, see Tables 4, 6 and 7. The obtained values were used to study the behavior of shifts in hyperfine intervals as a function of both temperature and $Z$, see Figs. 5, 6.
@@
The most important result of the research presented in this work is the discovery of the magic temperature at which the thermal frequency shift is zero [27]. For the most accurately measured frequencies in the hydrogen atom, as well as for some hydrogen-like ions, the results are presented in Table 5, see also Figs. 3, 4. It should be noted that there is a significant difference in the magic temperature for various hydrogen-like systems. For example, for doubly ionized lithium, the thermal shift of the two-photon 1s3s1s-3s transition frequency vanishes at a temperature significantly exceeding room temperature. At the same time, for lighter systems, magic wavelengths appear achievable in experiments involving two-photon transitions 1s2s1s-2s, 1s3s1s-3s (except for the helium ion in the latter case). It should also be noted that the analysis, taking into account the hyperfine structure of atomic levels, revealed that the latter is insignificant for determining $T_{m}$. This can be seen directly from Table 7: the thermal shifts to the hyperfine intervals of the ground and $2s$ states were always positive, differing by approximately an order of magnitude and at the same time negligibly small. For a hydrogen atom and the most accurately measured transition frequencies of 1s2s1s-2s and 1s3s1s-3s [46], the values obtained are $T_{m}=65.9$ K and $T_{m}=29.7$ K, respectively.
[…]
black It is well established that a "magic wavelength" can be used, at which the shifts of the atomic energy levels are equal and thus cancel in the definition of the transition frequency [48]. A similar concept can be applied to the thermal shift, i.e., when the levels between which the transition frequency is measured are shifted by the same amount. According to the obtained results, establishing and controlling this $T_{m}$ (magic temperature) appears to be feasible, see Table 5. In our view, stabilizing the temperature of the ambient radiation (from the chamber walls) in combination with employing the magic wavelength technique to cancel the level shift induced by the trapping laser would allow, at least to the lowest order, the influence of thermal radiation to be avoided. A more complex (as it necessitates the corresponding measurements) but realistic alternative appears to be extrapolating the data obtained at different temperatures to the point where the total shift vanishes.
[…]
black The generalization of the obtained results to multi-electron systems can be achieved in several ways. One approach is based on a quantum mechanical treatment (now commonly accepted) within second-order perturbation theory, which fully corresponds to the non-relativistic limit of the expressions used here; see also the discussion in [26]. In this case, one typically calculates the dipole polarizability of a multi-electron atom and the corresponding dynamic corrections arising from the series expansion of the dynamic polarizability at the low frequency of the thermal radiation. Another possible way is the use of semi-empirical methods, which allow for an efficient description of multi-electron atoms within a single-electron approximation, see, e.g., [40, 38]. In this case, the computational method presented in this work can be employed. Finally, the most complex but rigorously consistent method involves a generalization of the one-electron diagrams to the multi-electron case, see, for instance, [25]. We leave the latter two possibilities for future investigations.